# Proton Radiation Damage and Annealing of COSI p-type Cross-strip HPGe Detectors

### v1

@@ 4.3 Post Room Temperature Cycling
Previous studies report detrimental effects on the spectral resolution of radiation-damaged detectors cycled to room temperature [15, 13, 22, 14]. To understand the extent of this damage, we carried out our initial anneal by bringing the detector to C for 336 hours. The photopeak we measured after this anneal, shown in Figure 3, appears nearly bimodal, making it difficult to characterize accurately with one FWHM measurement. At this point, the detector was rendered unusable as a spectrometer.
Furthermore, the line profile is unlike any of the profiles we observed after solely primary damage. As expected, a small peak appears at the expected 0.662 MeV, however, the physical origin of the more prominent peak is unclear. It is downshifted to approximately 0.520 MeV and merging with the Compton backscatter continuum we expect from Cs-137. This shift is due to extreme trapping, demonstrating that the detector had a high trap density. However, in our charge transport models simulated with high trap densities/ low trapping products, we fail to observe this artifact. This suggests that the nature of the additional trapping caused by annealing at room temperature is separate from the trapping we observed from proton irradiation and that which is modeled in our numerical simulations. The distinction between these types of traps is important to maintaining the resolution of our GeDs and warrants more extensive study beyond the scope of this work.
@@ 4.4 High-temperature Anneals
After the first two anneals at C, summing to 72 hours, we observed FWHMh reduced to 21.11 keV for hole-collecting strips. The photopeaks measured after each of the five C anneals are plotted in Figure 4 and the repair of FWHMd with time annealing at C is marked by blue datapoints and fit to an exponential function plotted in Figure 5. The time constant for this repair is hours. This fit projects a detector spectral resolution, as measured by FWHM, will be repaired within 10 of its preradiation value in hours annealing at C. When we fit this decay we assume that there is no permanent damage induced and all FWHMd can be repaired by high-temperature anneal. Additionally we assume only the population of primary defects is present after the C anneals. We justify this assumption in Section 6.1.
We use our simulation results (section 3.3) to relate Cs-137 FWHM to the trapping product (Table 2) in order to track the way in which trap density changes with time annealing at C. We plot the results in Figure 6. The exponential decay fitted to this curve has a time constant of hours, a rate of repair consistent with that of FWHMd.
5 Conclusions
We irradiated COSI GeDs with the maximum fluence expected over an extended year mission, and induced additional damage by cycling the detector to room temperature, increasing the FWHM by 60 time is preradiation value. Through high-temperature annealing, we are able to repair the spectral resolution to within 16 of its preradiation value. This near full recovery of detector resolution demonstrates that high-temperature annealing can repair even the most heavily damaged detectors. Furthermore, our measurements enable us to characterize a repair timescale for high-energy proton damage.
The high proton fluences used in this study, along with secondary damage from room temperature cycling, significantly overestimate the damage that COSI will receive during its primary mission. Guided by the linear relationship between fluence and measured FWHM demonstrated in this study, we calculate a linear regression and obtain the projected spectral resolution at the end of COSI’s primary 2-year mission. We apply the repair timescale derived in this work to estimate the anneal time needed to repair this resolution to within 10 of our current preradiation value. From these results, we anticipate that COSI spectral resolution can be repaired by annealing 100∘C in hours, on the scale of two weeks. Confirming this repair time by conducting irradiation and annealing studies at lower proton fluences will be the subject of future work.
@@ 6 Discussion
Our FWHM measurements and spectra collected after irradiation and room temperature cycling (detailed in Sections 4.2 and 4.3, respectively) indicate that there are two distinct components to induced trapping, which we have identified as the primary and secondary components. This observation is confirmed by the study of secondary defects in Fourches (1991a) and Fourches (1991b). Our goal with these first two anneals was to determine if primary damage could be annealed at lower temperatures than previously used (C) on a reasonable timescale. Upon inspection, these first two anneals appear to have rapidly repaired secondary damage caused by the cycling to room temperature with little impact on primary damage.
Annealing studies completed on damaged p-type germanium have shown that the repair time of high-temperature annealing increases exponentially with decreasing temperature [18, 23]. In these studies, a 20∘C temperature decrease corresponds roughly to a repair time which is five times longer [18]. To repair primary damage, previous studies report annealing times on the scale of hundreds of hours for degradations 10 -100 times larger than the target resolution, such as that which is induced in this work [22, 15]. We similarly observe repair times on the scale of hundreds of hours at C as demonstrated in Section 4.4. This would imply that repair times at C must be five times longer, on the scale of thousands of hours. However, in our work, we initially see a relatively rapid improvement in resolution when annealing at C, on the timescale of tens of hours, but saturating at a resolution much degraded compared to the undamaged detector. This motivates us to assume that the secondary defects present prior to 80∘C anneals must repair on a relatively shorter timescale ( 72 hours) at lower temperatures, while the primary defects require higher temperatures (100∘C) and significantly longer anneal times.
Tracking only resolution attributed to external damage, we plot our measurements for FWHMd for hole-collecting strips across C anneals as blue data points in Figure 7 and attempt to fit decays of primary and secondary damage. We fit a double exponential function (FWHM) with two time constants referring to the timescales for primary and secondary annealing to the decay of FWHMd plotted against time spent annealing. Consistent with the discussion above, we assume that the portion of FWHMd attributed to primary defects must stay relatively constant from reaching room temperature to the end of C annealing (i.e 72 hours). The resulting decays of FWHMd, along with FWHMd,pri and FWHMd,sec, are plotted in Figure 7 with a time constant for the repair of secondary defects hours.
This indicates that secondary defects, while detrimental to spectral performance, repair at lower temperatures and significantly faster than primary defects. We conclude that C annealing procedures will not repair primary defects on a reasonable timescale for an instrument in space.
@@ 6.2 Formation of Secondary Defects
We find that, for hole-collecting strips, primary defects remaining after the room temperature cycling must account for 21 keV, a smaller amount of damage than we observe from the FWHM after the second round of radiation ( 50 keV). This suggests that a portion of the primary defects observed after radiation have been transformed into secondary defects when the detector was cycled to room temperature. This effect is observed in Fourches (1991b) which compares percentages of primary defects before and after room temperature cycling, and suggests that vacancies in disordered regions migrate out of these regions formed from primary damage and associate with each other to produce divacancies or clusters responsible for secondary damage.
Additional questions remain about the nature of traps from secondary damage. Previous literature argues that these are deep traps in order to explain their effect on resolution [12]. However, our measurements taken at separate detector depths after room temperature cycling diverge from charge trapping models outlined in [17], perhaps indicating detrapping effects. This would suggest, contrary to the conclusions of this earlier work, that secondary damage produces shallow traps which are repaired on a shorter timescale and at lower temperatures compared to primary damage, as demonstrated in Section 6.1. Answering these questions is beyond the scope of this work, but could be key to maintaining the spectral resolution of GeDs operating in the space environment.
@@ 6.3 Permanent Damage
The remaining damage-induced spectral resolution after 550 hours of 100∘C annealing may be evidence of permanent damage. We spent 192 hours annealing at 100∘C between anneals 6 and 7 and observe no improvement in detector resolution. The measured FWHM for hole-collecting strips after anneals 6 and 7 remains 3 keV with FWHMd 1.5 keV. This suggests that there is some small component of damage which our high-temperature annealing cannot eliminate. It is unclear whether this remainder is sustained from irradiation or cycling to room temperature and if it could be repaired by longer or higher temperature annealing procedures. A future study, where irradiation and high-temperature annealing procedures are repeated without cycling the detector to room temperature, could answer these questions.

### v2

@@ 4.3 Post Room Temperature Cycling
Previous studies report detrimental effects on the spectral resolution of radiation-damaged detectors cycled to room temperature [16, 14, 23, 15]. To understand the extent of this damage, we carried out our initial anneal by bringing the detector to C for 336 hours. The photopeak we measured after this anneal, shown in Figure 5, appears nearly bimodal and the detector was rendered unusable as a spectrometer.
Furthermore, the line profile is unlike any of the profiles we observed after solely primary damage. As expected, a small peak appears at the expected 0.662 MeV, however, the physical origin of the more prominent peak is unclear. It is downshifted to approximately 0.520 MeV and merged with the Compton backscatter continuum we expect from Cs-137 around 0.48 MeV. This downward shift, which can also be observed in photopeaks prior to anneal 4 (Figure 6), is due to extreme trapping, demonstrating that the detector had a high trap density. However, in our charge transport models simulated with high trap densities/ low trapping products, we fail to observe this artifact. As demonstrated in Figure 7, simulated photopeaks with lower trapping products have FWHM measurements similar to those of their experimental counterparts. However, their line profiles differ as simulated distributions drop off steeply at lower energies instead of containing a distinct downshifted peak. This suggests that the nature of the additional trapping caused by annealing at room temperature is separate from the trapping we observed from proton irradiation and that which is modeled in our numerical simulations. The distinction between these types of traps is important to maintaining the resolution of our GeDs and warrants more extensive study beyond the scope of this work.
@@ 4.4 High-temperature Anneals
After the first two anneals at C, summing to 72 hours, we observed FWHM reduced to 37.2 keV for hole-collecting strips. The photopeaks measured after each of the five C anneals are plotted in Figure 6 and the repair of FWHM with time annealing at C is marked by blue data points in Figure 8(a). We characterize a time scale for repair of primary damage at C by fitting these data points to an exponential function plotted in red. Statistical errors on these data points are small ( 0.2 keV FWHM) compared to their scatter, so we assume the uncertainties on the fit parameters are dominated by systematic uncertainties, which were adjusted to achieve a reduced chi-squared value of unity. We omit data points from anneals 2 and 3 from this fit as their line profiles, pictured in Figure 7, show evidence of remaining secondary damage. The time constant best fit to the repair of primary damage is hours. To account for potential permanent damage which cannot be repaired by annealing at 100∘C, we included a constant term which is best fit at keV. We apply the above time scale to the primary damage induced in this study and calculate a repair within 1% of FWHM = 2.4 keV in 49030 hours of annealing at C.
We use our simulation results (Section 3.3) to relate Cs-137 FWHM to the trapping product (Table 1) in order to track the way in which trap density changes with time annealing at C. We plot the results in Figure 8(b). The exponential decay fitted to these data, again omitting measurements from anneals 2 and 3, has a time constant of hours, a rate of repair consistent with that of FWHM.
@@ 5 Discussion
Our FWHM measurements and spectra collected after irradiation and then after room temperature cycling (detailed in Sections 4.2 and 4.3, respectively) indicate that there are two distinct components to induced trapping, which we have identified as the primary and secondary components. This observation is confirmed by the study of secondary defects in Fourches et al. [14] and Fourches et al. [13]. Each type of defect will repair, under high temperature anneal, on a different timescale. Thus we model the repair of defects with time spent annealing as a double exponential function:
We attempt to determine the two time constants at 80∘C for primary and secondary damage repair, and .
We rely on our time constant derived for C annealing in Section 4.4 to determine a repair time for primary damage at 80∘C. Annealing studies completed on radiation damaged p-type germanium have shown that the repair time of high-temperature annealing increases exponentially with decreasing temperature [19, 24]. In these studies, a 20∘C temperature decrease corresponds roughly to a repair time that is five times longer [19]. Therefore, 590 hours.
We argue in Section 4.4 that the time constant obtained for annealing at C tracks the rate of repair for primary defects alone. Comparison of experimental and simulated line profiles in Figure 7 reveals bimodal line profiles for anneals 2 and 3 characteristic of the distribution produced by room temperature cycling. By comparison, anneals 4-7 have single-peaked distributions which are in alignment with our simulations that model primary damage. This motivates us to omit measurements from anneals 2 and 3 from the fitted decay to obtain a repair of primary damage at C. Furthermore the time scale we obtain from this fit (118 hours) is consistent with previous studies that report annealing times for repair of primary damage at C on the scale of 100 hours [23, 16].
To derive a rate for the repair of secondary damage at C, we fit our measurements for FWHM for hole-collecting strips across C anneals to Equation 4. We fix 59060 hours following the discussion above and A = 251.5 keV in accordance with the amplitude of our fit in Figure 8(a). We plot data points in blue in Figure 9 and attempt to fit decay of secondary damage, B (the amplitude of FWHM), and a constant attributed to damage that cannot be repaired by C annealing. The resulting decay of FWHM, together with FWHM and FWHM, is plotted in Figure 9. We obtain a time constant for the repair of secondary defects hours.
This indicates that secondary defects, while detrimental to spectral performance, repair at lower temperatures and significantly faster than primary defects. This analysis appears consistent with the assumption that C annealing procedures will not repair primary defects on a reasonable timescale for an instrument in space.
@@ 5.2 Formation of Secondary Defects
Additional questions remain about the nature of traps from secondary damage. Previous literature argues that these are deep traps in order to explain their effect on resolution [13]. However, the spectral peaks shifts inferred at separate detector depths from data collected after room temperature cycling diverge from charge trapping models outlined in Boggs et al. [18], perhaps indicating detrapping effects. This would suggest, contrary to the conclusions of this earlier work, that secondary damage produces shallow traps which are repaired on a shorter timescale and at lower temperatures compared to primary damage, as demonstrated in Section 5.1. Answering these questions is beyond the scope of this work, but could be key to maintaining the spectral resolution of GeDs operating in the space environment.
@@ 5.3 Permanent Damage
The remaining damage-induced spectral resolution after 550 hours of 100∘C annealing may be evidence of permanent damage. We spent 192 hours annealing at 100∘C between anneals 6 and 7 and observe only 0.3 keV improvement in detector resolution. The measured FWHM for hole-collecting strips after anneal 7 remains 4.08 keV with FWHM 2.79 keV. This suggests that there is some small component of damage which our high-temperature annealing at 100∘C cannot eliminate. It is unclear whether this remainder is sustained from irradiation or cycling to room temperature and if it could be repaired by longer or higher temperature annealing procedures. A future study, in which irradiation and high-temperature annealing procedures are repeated without cycling the detector to room temperature, could answer these questions.
6 Conclusions
We irradiated COSI GeDs with the maximum fluence expected over an extended 10 year mission, and induced additional damage by cycling the detector to room temperature, increasing the FWHM by 38 times is preradiation value. Through high-temperature annealing we are able to repair the spectral resolution to within 37 of its preradiation value, repairing to 4.08 keV compared to the original 2.98 keV. This notable recovery of detector resolution demonstrates that high-temperature annealing can repair even the most heavily damaged detectors. Furthermore, our measurements enable us to characterize a repair timescale for high-energy proton damage.
The high proton fluences used in this study, along with secondary damage from room temperature cycling, significantly overestimate the damage that COSI will receive during its primary mission. Guided by the linear relationship between fluence and measured FWHM demonstrated in Section 4.2, we calculate a linear regression and obtain the projected spectral resolution at the end of COSI’s primary 2-year mission (FWHM 8.8 keV ). We apply the repair time scale derived in this work to estimate the anneal time needed to repair this resolution within 1% of damage unrepairable at 100∘C (FWHM = 2.4 keV). From these results, we anticipate that COSI spectral resolution can be repaired by annealing 100∘C in hours, on the scale of two weeks. Confirming this repair time by conducting irradiation and annealing studies at lower proton fluences will be the subject of future work.